The Skysong Corporation issued 10-year, $5,270,000 par, 6% callable convertible subordinated debentures on January 2, 2017. The bonds have a par value of $1,000, with interest payable annually. The current conversion ratio is 13:1, and in 2 years it will increase to 19:1. At the date of issue, the bonds were sold at 96. Bond discount is amortized on a straight-line basis. Skysong’s effective tax was 35%. Net income in 2017 was $8,800,000, and the company had 1,985,000 shares outstanding during the entire year.

(a) Compute both basic and diluted earnings per share. (Round answers to 2 decimal places, e.g. $2.55.)

Basic earnings per share $

Diluted earnings per share $
Answer #1
Answer
Step-1 Basic earnings per share
Net income / Average number of shares outstanding
.$7=88,00,000 / 1985000
$4.4
Step-2 Compute annual interest expense on the bonds.
Annual interest expense on the bonds
Cash interest paid + Discount on bonds amortized annually
($5270,000 x 6%) + ($5270,000 x 4% x 1/10)
$337280
Step-3 Compute the amount of after tax savings if the bonds are converted in to shares of common stock.
Savings = Interest expense x (1 - tax rate) = $337280 x (1 - 0.35) = $219232
Step-4 Compute net income if the bonds are converted in to common stock.
Net income after conversion
Net income before conversion + Savings in interest expense if bonds are converted
$8800,000 + $219232
$9019232
Step-5 Compute the number of shares to be issued to convert the bonds after two years at a conversion ratio of 19:1
Number of shares to be issued = (5270,000/$1,000) x 19 = 100,130
Compute the total number of shares outstanding after conversion of bonds.
Total number of shares after conversion of bonds = 1985000 + 100130 = 2085130
Step-6 Compute diluted earnings per share.
Diluted earnings per share
Net income after conversion / number of shares outstanding after conversion
$9019232 / 2085130
$4.33